OpenStudy (anonymous):

Is the sequence geometric? If so, identify the common ratio. 6, 12, 24, 48, ... (1 point)yes; 2 yes, –2 yes; 4 no

OpenStudy (anonymous):

@amistre64 @Hero

OpenStudy (anonymous):

yes;2 6*2=12 12*2=24 24*2=48